# A theory of quasiballistic spin transport

### v1

@@ I Introduction
Thus, unlike in continuous systems, faster-than-diffusive transport in lattice models is generally linked to the presence of integrability [7, 4]. As one perturbs away from a given integrable point, such faster-than-diffusive dynamics usually gives way to normal diffusion. However, the relaxation timescales controlling this crossover to normal diffusion can be extremely long, even when the system under consideration is chaotic in all other respects [8, 9, 10, 11, 12, 13, 14, 15].
In particular, several recent works have observed numerical evidence for long-lived and transiently ballistic (or “quasiballistic”) dynamics of conserved charges in systems that appear to be far from integrability. For example, spin transport in power-law-interacting spin-$1/2$ XXZ chains [13, 14] and charge transport in interacting fermionic lattice models in any spatial dimension [15] can exhibit this phenomenon. Both sets of work focus on infinite temperature, where such behaviour is intuitively least likely to occur. In the fermionic case, this physics was explained theoretically by computing an effective lifetime for the charge current, which could be made arbitrarily small as a function of the interaction range. Meanwhile, no detailed theoretical explanation has been proposed for quasiballistic spin transport and providing such an explanation is the primary goal of this work.
@@
We emphasize that Hamiltonians of this type are not merely a theoretical curiosity. Recent advances in quantum devices have renewed interest in the dynamics of long-range interacting spin chains [16, 17, 18, 19, 20]. Importantly, these platforms inherently favour the study of real-time dynamics. Moreover, their relatively high effective temperatures, which can be undesirable when it comes to exploring traditional questions about ground-state physics, are ideal for studying the far-from-equilibrium transport properties of interest in this work. There now exist several concrete proposals [21, 22, 23] for how to realize precisely the power-law interacting XXZ Hamiltonians studied here using trapped ions and a recent experiment [24] probing their dynamics. It should, however, be borne in mind that relatively small power-law exponents $0\leq\alpha\leq 3$ are most physically natural in these platforms [25], while the technique in our paper is simplest to apply for $\alpha>3/2$.
@@
In this work, we pursue a less ambitious yet analytically tractable version of the path sketched above. Rather than ruling out ballistic spin transport fully for $\alpha\neq 2$, we rule out the possibility of ballistic spin transport protected by a strictly two-body conservation law, as it is at the Haldane-Shastry [26, 27, 4] and XX points. Similarly, rather than attempting to evaluate the spin diffusion constant from the Kubo formula, we estimate the lifetime of the spin current directly from its instantaneous rate of change, a strategy that has proved effective in fermionic systems [15] and is reasonable in the absence of ballistic transport. We conclude by discussing the scope of our approach and the connection between integrability and transport more broadly. In particular, we prove that the only spin-$1/2$ chains of the form Eq. (1) with an exactly conserved spin current are Haldane-Shastry and XX chains (Theorem 1 of Appendix A.1).
@@ II Decay rate of the spin current
where $a(r)$ and $b(r)$ are arbitrary even functions and it will be convenient  [^1] The analysis in our paper could in principle be extended to longer interaction ranges using e.g. a suitable variant of Kac normalization, but we will not pursue this direction here. to assume that $a(r),\,b(r)=o(r^{-3/2})$ as $r\to\infty$. We additionally assume that $a$ is non-zero (i.e. $a(r)\neq 0$ for some integer $r\neq 0$) to exclude Ising models without spin transport.
@@
For generic, chaotic, inversion-symmetric Hamiltonians we expect diffusive spin transport above zero temperature (see Appendix A.2 for details). In such systems, a reasonable estimate for the infinite-temperature decay rate of the spin current can be obtained as follows [15].
@@ III.3 Optimal relaxation rate and perturbative regime
While this is perhaps surprising for intermediate values of $\alpha$, it is expected that as $\alpha\to\infty$, power-law exchange interactions will merely yield a small (integrability-breaking) next-nearest-neighbour “hopping” correction $a(2)=2^{-\alpha}$ to the integrable XX point [13]. Moreover, for anisotropies $\Delta$ that are exponentially small in $\alpha$, the nearest-neighbour Ising “interaction” term $b(1)=\Delta$ will also be a small (integrability-preserving) perturbation to the XX point. This raises the questions of which small perturbation bottlenecks our estimate of the relaxation rate of the spin current for $\alpha\gg 1$ – on physical grounds, this should be the integrability-breaking hopping term – and how far our estimate for the optimal lifetime $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ improves upon the lifetime $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}_{\mathrm{num}}(\alpha))$ extrapolated from the previous numerical observations [13, 14] in this perturbative regime.
@@ IV Discussion
While we cannot exclude this possibility based on available theoretical techniques, we prove a partial result justifying our approach in Appendix A.2. Specifically, we show that for the models studied in this paper, there is no two-body conserved charge that could give rise to a non-zero Mazur bound for $\Delta\neq 0,\pm 1$. Unfortunately, this does not eliminate the possibility of quasilocal, many-body conserved charges that protect the spin current, as in the nearest-neighbour XXZ chain [31, 34]. At the same time, a systematic construction of quasilocal conserved charges for an arbitrary lattice model without prior knowledge of its underlying integrable structure is an extremely challenging problem whose solution would have immediate implications for various important open problems in many-body physics, such as the rigorous justification of many-body localization [35, 36, 37, 38].
@@
A corollary of our analysis is a simplified integrability-transport theorem (see Theorem 1 of Appendix A.1): for spin-$1/2$ chains of the form Eq. (1), we show that spin transport is only perfectly ballistic if the model is equivalent to either a nearest-neighbour XX chain or a Haldane-Shastry chain. Extending this result to short-range interactions, imperfect ballistic transport and quasilocal conserved charges, along the lines of recent rigorous results on non-integrability of certain spin chains [39, 40, 41], seems desirable for understanding the relationship between integrability and transport more generally.
@@ V Acknowledgments
We thank D. A. Huse for numerous helpful discussions and collaborations on related work and M. Mierzejewski for drawing quasiballistic spin transport to our attention. H. H. is supported by NSF QLCI grant OMA-2120757. W. W. H. is supported by the National Research Foundation (NRF), Singapore, through the NRF Felllowship NRF-NRFF15-2023-0008, and through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009).

### v2

@@ I Introduction
Thus, unlike in continuous systems, faster-than-diffusive transport in lattice models is generally linked to the presence of integrability [7, 4]. As one perturbs away from a given integrable point, such faster-than-diffusive dynamics usually gives way to normal diffusion. However, the relaxation timescales controlling this crossover to normal diffusion can be extremely long, even when the system under consideration is chaotic in all other respects [8, 9, 10, 11, 12, 13, 14, 15, 16].
In particular, several recent works have observed numerical evidence for long-lived and transiently ballistic (or “quasiballistic”) dynamics of conserved charges in systems that appear to be far from integrability. For example, spin transport in power-law-interacting spin-$1/2$ XXZ chains [14, 15] (see also [8]) and charge transport in interacting fermionic lattice models in any spatial dimension [16] can exhibit this phenomenon. Both sets of work focus on infinite temperature, where such behaviour is intuitively least likely to occur. In the fermionic case, this physics was explained theoretically by computing an effective lifetime for the charge current, which could be made arbitrarily small as a function of the interaction range. Meanwhile, no detailed theoretical explanation has been proposed for quasiballistic spin transport in general and providing such an explanation is the primary goal of this work.
@@
We emphasize that Hamiltonians of this type are not merely a theoretical curiosity. Recent advances in quantum devices have renewed interest in the dynamics of long-range interacting spin chains [17, 18, 19, 20, 21]. Importantly, these platforms inherently favour the study of real-time dynamics. Moreover, their relatively high effective temperatures, which can be undesirable when it comes to exploring traditional questions about ground-state physics, are ideal for studying the far-from-equilibrium transport properties of interest in this work. There now exist several concrete proposals [22, 23, 24] for how to realize precisely the power-law interacting XXZ Hamiltonians studied here using trapped ions and a recent experiment [25] probing their dynamics. It should, however, be borne in mind that relatively small power-law exponents $0\leq\alpha\leq 3$ are most physically natural in these platforms [26], while the discussion in our paper applies to $\alpha>3/2$.
@@
In this work, we pursue a less ambitious yet analytically tractable version of the path sketched above. Rather than ruling out ballistic spin transport fully for $\alpha\neq 2$, we rule out the possibility of ballistic spin transport protected by a strictly two-body conservation law, as it is at the Haldane-Shastry [27, 28, 4] and XX points. Similarly, rather than attempting to evaluate the spin diffusion constant from the Kubo formula, we estimate the lifetime of the spin current directly from its instantaneous rate of change, a strategy that has proved effective in fermionic systems [16] (see also [29]) and is reasonable in the absence of ballistic transport. We conclude by discussing the scope of our approach and the connection between integrability and transport more broadly. In particular, we prove that the only spin-$1/2$ chains of the form Eq. (1) with an exactly conserved spin current are Haldane-Shastry and XX chains (Theorem 1 of Appendix A.1).
@@ II Decay rate of the spin current
where $a(r)$ and $b(r)$ are arbitrary even functions and it will be convenient  [^1] For power-law exponents 1/$21/2<\alpha\leq 3/$ 2, there is evidence [18, 20] that transport in models of this type crosses over to superdiffusion, which in turn gives way to unphysical instantaneous relaxation for $\alpha<1/2$. Indeed, the analysis in our paper is complicated for $\alpha\leq 3/2$ by the failure of Eq. (7) to converge. Restoring a finite-size cutoff in Eq. (8) and performing a numerical finite-size scaling analysis, not shown, we see behaviour consistent with the qualitative picture proposed in Ref. [18] and therefore with the absence of quasiballistic transport for $\alpha\leq 3/2$. to assume that $a(r),\,b(r)=o(r^{-3/2})$ as $r\to\infty$. We additionally assume that $a$ is non-zero (i.e. $a(r)\neq 0$ for some integer $r\neq 0$) to exclude Ising models without spin transport.
@@
For generic, chaotic, inversion-symmetric Hamiltonians we expect diffusive spin transport above zero temperature (see Appendix A.2 for details). In such systems, a reasonable estimate for the infinite-temperature decay rate of the spin current can be obtained as follows [16, 29].
@@ III.3 Optimal relaxation rate and perturbative regime
While this is perhaps surprising for intermediate values of $\alpha$, it is expected that as $\alpha\to\infty$, power-law exchange interactions will merely yield a small (integrability-breaking) next-nearest-neighbour “hopping” correction $a(2)=2^{-\alpha}$ to the integrable XX point [14]. Moreover, for anisotropies $\Delta$ that are exponentially small in $\alpha$, the nearest-neighbour Ising “interaction” term $b(1)=\Delta$ will also be a small (integrability-preserving) perturbation to the XX point. This raises the questions of which small perturbation bottlenecks our estimate of the relaxation rate of the spin current for $\alpha\gg 1$ – on physical grounds, this should be the integrability-breaking hopping term – and how far our estimate for the optimal lifetime $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ improves upon the lifetime $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}_{\mathrm{num}}(\alpha))$ extrapolated from the previous numerical observations [14, 15] in this perturbative regime. (We note that the emergence of quasiballistic spin transport from these specific perturbations to the XX model was studied previously using different techniques [8].)
@@ IV Discussion
While we cannot exclude this possibility based on available theoretical techniques, we prove a partial result justifying our approach in Appendix A.2. Specifically, we show that for the models studied in this paper, there is no two-body conserved charge that could give rise to a non-zero Mazur bound for $\Delta\neq 0,\pm 1$. Unfortunately, this does not eliminate the possibility of quasilocal, many-body conserved charges that protect the spin current, as in the nearest-neighbour XXZ chain [33, 36]. At the same time, a systematic construction of quasilocal or pseudolocal [36] conserved charges for an arbitrary lattice model without prior knowledge of its underlying integrable structure is an extremely challenging problem whose solution would have immediate implications for various important open problems in many-body physics, such as the rigorous justification of many-body localization [37, 38, 39, 40].
@@
A corollary of our analysis is a simplified integrability-transport theorem (see Theorem 1 of Appendix A.1): for spin-$1/2$ chains of the form Eq. (1), we show that spin transport is only perfectly ballistic if the model is equivalent to either a nearest-neighbour XX chain or a Haldane-Shastry chain. Extending this result to imperfect ballistic transport and pseudolocal conserved charges, along the lines of recent rigorous results on non-integrability of certain spin chains [41, 42, 43], seems desirable for understanding the relationship between integrability and transport more generally.
@@ V Acknowledgments
We thank D. A. Huse for numerous helpful discussions and collaborations on related work and M. Mierzejewski for drawing quasiballistic spin transport to our attention and comments on the manuscript. H. H. is supported by NSF QLCI grant OMA-2120757. W. W. H. is supported by the National Research Foundation (NRF), Singapore, through the NRF Felllowship NRF-NRFF15-2023-0008, and through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009).